OpenStudy (anonymous):

What is the slope and y-intercept for y=2x+3?

OpenStudy (harsha19111999):

compare the equation with y=mx+c

OpenStudy (harsha19111999):

m is the slope. m=2. therefore slope=2

OpenStudy (anonymous):

Okay so this?; y=(2)2+3 y=4+3 y=7? No, that isn't an option. I'm so confused...

OpenStudy (harsha19111999):

Wait where from did you get all other values? like x,c

OpenStudy (anonymous):

I have no idea. I'm just so confused with the whole thing.

OpenStudy (harsha19111999):

see you have the equation as y=2x+3. compare it with y=mx+c. Here m is the slope and c is the y intercept

OpenStudy (nincompoop):



OpenStudy (harsha19111999):

You asked for slope and y-intercept, so slope=2 and y intercept c=3

OpenStudy (harsha19111999):

Did u understand?

OpenStudy (nincompoop):

CLEARLY from the format, SLOPE is the value of m and the Y-INTERCEPT is the value of b

OpenStudy (anonymous):

Let's put it to you this way. fro the equation y = mx +c, m( the co-efficient of x) is your slope and c ( your constant is your y-intercept).

OpenStudy (solomonzelman):

dang it.... to answer your question you would need to know what a slope and y-intercept mean in general first. You apparently don't know....

OpenStudy (anonymous):

I know what they mean. Y-intercept is where the line crosses the y axis. Slope is like an elevated line on one side.

OpenStudy (anonymous):

yes.. so where are you getting stuck?